OpenStudy (anonymous):

find the derivative of f(x)= x(x^2-1) / (x+3). Simplify your answer

OpenStudy (dumbcow):

i think quotient rule here f = x^3 - x f' = 3x^2 - 1 g = x+3 g' = 1 \[(\frac{f}{g})' = \frac{f'g - fg'}{g^{2}}\]

OpenStudy (anonymous):

so that would be 3x^2-1(x+3) / 1^2

OpenStudy (anonymous):

what would I do from there?

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

because wolframalpha says the derivative is f'(x) = 2x^3+9x^2-3 / (x+3 )^2 but I dont know how it got to thtat

OpenStudy (anonymous):

please help me?

OpenStudy (amtran_bus):

@dumbcow